The CDF Experiment

Study of p p interactions at √s = 1.8 TeV

• 25 institutions

• 400+ physicists

IPP Group consists of

McGill University  University of Toronto
Faculty (1.4)  Ken Ragan (0.7)  Pekka Sinervo (0.7)
RA (2.0)  Kurt Biery (1.0)  Andreas Hölscher (0.5)

Klaus Strahl (0.5)

Grad Students (5.0) Yunling Ye George Sganos
Kostas Kordas  Andreas Warburton

Hyunsoo Kim

Group is in the middle of data-collecting for Run I

• Run IA (9/92 through 5/93) 18 pb−1⇒∼ 25 invpb

• Run IB (10/93 through 6/94) ⇒∼ 75 invpb

Physics Programme

Physics goals of Run I

• Discover top quark if ∼ Mtop < ∼ 140 GeV/c2

• Bottom Quark Physics – Search for new states (Bs and Λb)

– Measure lifetimes and decay widths of b hadrons – Search for rare b decay modes

• Electroweak Measurements – Production properties of W ±’s and Z∘’s

– Measure MW to precision < ∼ 0.15 GeV/c2

– Search for new currents (W ′ and Z′)

• QCD Studies

• Search for Exotica

Top Quark Search

CDF currently has proven capability for top search

• Current best preliminary limit is from CDF ⇒ 40 < Mtop < 108 GeV/c2 (95% CL)

• Includes all decay channels with e±’s and μ±’s

• Have unique vertex B tagging

Lepton + Jets Channel is an IPP group focus

• electron + neutrino ID eliminates QCD background

• Remaining background from W + jets production

• Removed with B tagging

Data from current run being analyzed

• Any possible excess consistent with Mtop ∼ 160 GeV/c2 • Signal-to-noise is extraordinarily difficult problem

Bottom Quark Physics

Bottom hadron spectroscopy

• Bs and Λb search

• Measure lifetimes with SVX detector

• Measure branching ratios

• Bs − Bd mixing

Bottom Quark Production Cross Section

• Measured σb using – b → e/μ+ νeX

– B → J/ψK±

– Inclusive e± sample dominated by b production • Measure σc using χ → J/ψγ mode

CDF Detector Upgrades

Upgrades in place for Run IA

• Central muon upgrade and extension to |η| < 1.0

• SVX Detector (4 layers of 55 and 65 μ pitch Si strips)

• DAQ improvements

• L3 SGI farm (⇒> 1200 MIPS)

• Offline and analysis software ported to UNIX

Upgrades for Run IB

• Install first part of DAQ upgrade

• SVX II detector (complete replacement for SVX)

• Trigger improvements and L3 upgrade (×2)

IPP Contributions

Operating Support (to date or committed)

• Limited support from NSERC ($140K/year for 91/92/93)

• Startup support from McGill and Toronto

• NSERC support of $180K/year for 93/94/95

Equipment Support (to date or committed)

• $150K from startup funding

• $130K for 92/93 DAQ Upgrade

• Computing upgrade for McGill (91/92) ($25K)

• Computing upgrade for Toronto (93/94) ($30K)

Detailed Breakdown of IPP Effort

Hardware Responsibilities:

• Current DAQ System support (Biery/Sinervo)

• Run IB DAQ Upgrade (Hölscher/Biery/Strahl/Ragan/Kim)

• Offline software support (Sinervo/Warburton/Ragan)

• Detector Simulation (Sganos/Ragan/Kordas/Ye)

Physics Analyses:

• Bs search (Sganos/Strahl/Hölscher)

• χc production (Ye)

• B → ψ′X decays (Warburton)

• W ′ search (Kim)

• Top search (Sinervo)

Long Range Plans for Tevatron

Run I will see ∼ 100 pb^{−1} delivered to CDF & D0

• Improved Tevatron luminosity of ∼ 7 × 1030 cm−2s−1 • Achieved with electrostatic seperators/improved p̄ stacking • New linac

Run II will use Main Injector

• Increase collider aperture for proton and p̄ injection

• Cool collider to 3.9∘K ⇒ √s = 2.15 TeV

• Increase injected beam intensities

• Goal is Lmax ∼ 5 × 1031 cm−2s−1 and 500 pb−1

• Expected completion is fall 1996

Funding Profile

Goal is to complete Physics Analysis of Run 1A/IB

• Data for Run IB collected by summer 1994

• Most physics analysis complete by end of 1995

• Manpower would remain approximately constant

Have to consider a number of scenarios

1. CDF programme winds down as analyses complete

• Transition period as student’s work completes

• Funding begins to fall off in 96/97

• Group focuses on new initiatives

2. Continue CDF collaboration through Run II

• Requires funding to complete DAQ upgrade ($150K)

• Extend IPP involvement in CDF through 98/99

Why Consider Run II?

Exciting Physics Opportunities

• Only place to study top quark system?

• CP violation?

• New physics that is uncovered in Run I?

Schedule Slippage in New Initiatives

• SSC or LHC turn-on after 2000?

• B factory not constructed?

Estimated IPP Contributions:

• Operating support ($250K/year through 97/98)

• Equipment for completion of DAQ upgrade (∼ 150 − 200K)

Summary

• CDF will continue data-taking through till 1994 • IPP should continue to be involved through 95/96

– Earlier cut-off not a reasonable option

• Possibility of participating in Run II